OpenStudy (anonymous):

Express the radical in simplest form:square root of 45(x) A:9x square root of 5 B:3 square root 5x C:5 square root 3x

OpenStudy (lgbasallote):

\[\Large \sqrt{45x} = \sqrt{45 \times x} = (\sqrt{45}) \times (\sqrt x) = (\sqrt{9 \times 5}) \times (\sqrt x) = (\sqrt 9) \times (\sqrt 5) \times (\sqrt x)\] \(\sqrt 9\) is a perfect square isnt it?

OpenStudy (anonymous):

so whats the answer?

OpenStudy (lgbasallote):

what is the square root of 9?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

A,B, OR C?